1. ABC Corp. has an ROE (return on reinvested earnings) of 20% and a dividend payout ratio of 40%. The next annual earnings are expected to be $3 per share (that is, EPS in year 1 is $3.00). The firm's required return on the stock is 17%. The value of the stock today is $____________.

2. Company A  just paid a $1.00 dividend per share and its future dividends are expected to grow at an annual rate of 6% for the foreseeable future. The beta of company A's stock is 1.25, the risk-free rate of return is 4% and the expected return on the market portfolio is 10.4%. The value of the stock today is $___________.
Answer #1

r CROE) = 20% (0) EPS = $3 (E) durdend payout = 40% Re = 17% (Re) (D) or 40% of 3 = 12 By, Walter model ie. P = D + (Re) (E-DGwen : Do =$1 .:D = Do (1+g) - g = 6% perpeluăl = 1(1+0.06) = 1100 B = 1.25 rs = 4% Rm = 10.4% First by using CAPM model fundвч 9 amglaur mоuѕ mеdok at 9ария е . Re = D) +9 05. Po e Do Clta - Po - Re-g - о - ) - 06 0.12 - 0.06 = $19.61